OpenStudy (anonymous):

Find two positive integers greater than one whose product is 24,999,999 and whose positive difference is as large as possible. Now, I understand that the positive integers could be 124,999,99.5. It's the rest of the problem that confuses me.

OpenStudy (anonymous):

Oh-and medals for the correct answer!

OpenStudy (amistre64):

xy >k x-y = max ??

OpenStudy (amistre64):

either way you will have to solve a quadratic

OpenStudy (amistre64):

if xy = k, then we can sub in y = k/x f(x) = x - k/x is max when f' = 0

OpenStudy (anonymous):

What is a quadratic?

OpenStudy (anonymous):

I don't understand what your saying, sorry. :(

OpenStudy (amistre64):

its a poly of degree 2, but that might be a stray thought ...

OpenStudy (amistre64):

let y = k/x therefore: f(x) = x - k/x f' = 1 + k/x^2 = 0 x^2 + k = 0 x = sqrt(k) seems like a good bound

OpenStudy (amistre64):

sqrt(24999999) is about 4999.999.... so let x = 4999 and y= 24999999/x = 5001

OpenStudy (anonymous):

I have to eat dinner sorry...

OpenStudy (anonymous):

So the answer is 5001?

OpenStudy (amistre64):

positive difference is as large as possible i think i found the "as small as possible" instead

OpenStudy (amistre64):

1 and 24 999 999 might be the max positive difference

OpenStudy (amistre64):

Find two positive integers greater than one whose product is 24,999,999 and whose positive difference is as large as possible. since the product of 1 and 24 999 999 is equal to 24 999 999; and the difference between them is 24 999 998 i would say that you are looking for x=1 and y = something approaching infinity .....